OpenStudy (anonymous):

If the half-life of a 20.0 g sample is known to be 24 minutes, how long will it take for only 5.0 grams of the sample to remain?

OpenStudy (anonymous):

5g is 1/4 of 20g, so that's 2 half-lives 2x24 = 48 minutes

OpenStudy (dumbcow):

\[e^{-24k} = \frac{1}{2}\] solve for k 5/20 = 1/4 \[e^{-kt} = \frac{1}{4}\] solve for t

OpenStudy (anonymous):

48